OpenStudy (anonymous):

Find y': sin x / x^2. y' = (sin x)'*x^2-(sinx)(x^2)' / (x^2)^2 y' = cos x * x^2 - (sin x)(2x) / x^4 then what?

OpenStudy (anonymous):

quotient rule here \[(\frac{f}{g})'=\frac{f'g-g'f}{g^2}\]

OpenStudy (anonymous):

you looking for first derivative yes?

OpenStudy (anonymous):

you have it. factor out an x from the numerator and cancel with one of them in the denominator. that is it

OpenStudy (anonymous):

yes. sorry - i should have said in the beginning i was using the quotient rule

OpenStudy (anonymous):

get \[\frac{x \cos(x)-2\sin(x)}{x^3}\]

OpenStudy (anonymous):

ok. this is where I am stuck: how do you get x cos x? To the point, how does the x move in front of cos x? wouldn't it be just cos x once the x is factored out?

OpenStudy (anonymous):

your numerator was \[\cos(x)x^2-2x\sin(x)=x^2\cos(x)-2x\sin(x)=x(x\cos(x)-2\sin(x))\]

OpenStudy (anonymous):

just the commutative law and factoring. this allows you to cancel one of the "x"s from the denominator

OpenStudy (anonymous):

now i see my error: (sin x)' does not equal cos. it equals cos x.